THE SPECTRUM OF METHYL FORMATE IN THE THZ REGION

\underline{M. TUDORIE,} T. R. HUET^{a}, L. MARGULES, M. GOUBET, Laboratoire de Physique des Lasers, Atomes et Molécules (PhLAM) UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; O. PIRALI, P. ROY, Ligne AILES, synchrotron SOLEIL, LOrme des Merisiers, Saint Aubin, BP 48, 91192 Gif-sur-Yvette, France; V. V. ILYUSHIN, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; I. KLEINER, LISA, CNRS/Universités Paris 12 et 7, 61 Avenue du Général de Gaulle, 94010 Créteil, France.

The THz spectrum of methyl formate-HCOOCH_{3} is currently investigated. At first a multi-pass cell having an optical path of 150 m coupled to an internal source of the Fourier Transform spectrometer of the AILES beamline, synchrotron SOLEIL (France), was used to obtain the methyl formate THz spectrum. Preliminary assignments of the pure rotation spectrum up to 80 cm^{−1}, and of the very weak torsion band v_{t} = 1-0 around 130 cm^{−1} are carried out. The assignments are based on the rotation-torsion energy levels calculated using the RAM approach^{b}. The particular interest in Δv_{t} = 1 torsion-rotation band lies in the direct experimental determination of the barrier height V_{3}, which up to now was determined from pure rotational transitions only, and consequently in the contribution to the improvement of the global study of the rotational levels in the lowest torsional states of methyl formate. Secondly further measurements using the synchrotron radiation are planned. The latest results will be presented.

aThis work is supported by the ANR-08-BLAN-0054 contract

bIlyushin, Kryvda, Alekseev, J. Mol. Spectrosc. (2009), doi: 10.1016/j.jms.2009.01.016